Laser Subcommittee Meeting 2021-06-19
Attendees: BlakeS, NickB, JC, BenC, MichaelO'C, BruceC, FletcherB
Next meeting(s): Four weeks: regular third Saturday 14:00 ? OK
Promoted, extra attendees!
LG500 "Little Red"
cutting 3mm MDF in centre and corners of the bed at 40mm/s
KH7050 "Kellogg/Middle": operational, was out-of-order
cutting 3mm MDF in centre and corners of the bed at 20mm/s
X end-stop sensor failed 2021-05-22
"FOTEK PL-05N", NPN, 10-30VDC
brown: +V, blue: 0V, black: out
Levelled, shim placed under honeycomb: thanks Fletcher+Blake
Mirror clean, mirror 2 adjustment, beam was hitting head too high
LC1290 "Big Red"
cutting 3mm MDF in centre and corners of the bed at 50mm/s
Levelled 2021-06-18, thanks GlennM
Empty compressor tanks, water trap
The hour meter started from 1.6 hours, at 20171210 , 400-512 beam-active cutting hours per annum
488.9 100W installed 20181222
1515.6 SPT-TR130W tube installed